The number of molecules of ATP produced in the lipid metabolism of a molecule of palmitic acid is : (1) 56 (2) 36 (3) 130 (4) 86 

Grade:12th pass

1 Answers

Anjali Ahuja
8 years ago
Hello Student
Oxidation of of palmitic acid yields 7 NADH + 7 FADH2 + 8 acetyl-CoA in 7 cycles of mitochondrial beta oxidation.
1NADH produces3 ATP molecules and 1FADH2 produces 2ATP molecules. Every acetyl-CoA yields 3 NADH + 1 FADH2 + 1 GTP (=ATP) during Krebs cycle.

Now, 7 NADH + 7 FADH2 + 8 acetyl-CoA = 7*3 + 7*2 + 8 (3*3 + 1*2 + 1) = 131

Now we have 131 ATP molecules. However we have to used 2 ATP molecules for the initial activation of every fatty acid that is going to be oxidized in the mitochondria.
So it is 131-2 = 129 ATPs.
Hope everything is clear :)